cho sasaki demonstrates microwave ovens at the national home show she is paid $7 for each of the first 7 demonstations and $11 for each demonstration for each demonstration over 7 what is her commision for a day in which she makes 12 demonstrations

Answer by checkley79(3341)

7*7+(12-7)*11=49+5*11=49+55=$104 ANS.